OpenStudy (anonymous):

find the number of terms in AP in which the first tem =5 the common diffenrence=3 and last term = 83

OpenStudy (lilg132):

26 terms

OpenStudy (lilg132):

actually no 27 terms

OpenStudy (lilg132):

remember to click good answer

OpenStudy (anonymous):

hey pleae explain

OpenStudy (lilg132):

ok so we know the 1st term is 5 so we take away the last term from the first term 83 - 5 = 78 then to find the amount of terms you will need to divide the answer by 3 which will give you the number of terms with 1 taken out so 78 / 3 = 26 now remember we took 1 term out in the beginning so you add on 1 more term to 26 26 + 1 = 27 so all together there are 27 terms